1. What is the present value of $400, three years in the future if the interest rate is 1%?

2. X has a term deposit that pays 10% a year and its value after 2 years will be $2,500. What is the present value of X term deposit?

3. A loan Z $10,000 and a year later, Z pays A $10,400. If the inflation rate during that year is 1.5%, what is the real interest rate that Z is paying to A?

Thank you.
Answer #1

1. Present value = Future value x (1 + rate)^-n

= 400*(1 + 1%)^-3

= $388.24

2. Present value = Future value x (1 + rate)^-n

= 2500*(1+10%)^-2

= $2066.12

3. A loan Z $10,000 and a year later, Z pays A $10,400. This implies we have

Present value = Future value x (1 + rate)^-n

(10000/10400)^(-1) = 1 + Nominal rate of interest

Nominal rate of interest = 4%. Inflation rate during that year is 1.5%

Real interest rate = Nominal rate of interest - Inflation rate = 4 - 1.5 = 2.5%.

Hence, real interest rate is 2.5%